# Revitalizing Multivariate Time Series Forecasting: Learnable Decomposition with Inter-Series Dependencies and Intra-Series Variations Modeling

### v2

@@ Abstract
Predicting multivariate time series is crucial, demanding precise modeling of intricate patterns, including inter-series dependencies and intra-series variations. Distinctive trend characteristics in each time series pose challenges, and existing methods, relying on basic moving average kernels, may struggle with the non-linear structure and complex trends in real-world data. Given that, we introduce a learnable decomposition strategy to capture dynamic trend information more reasonably. Additionally, we propose a dual attention module tailored to capture inter-series dependencies and intra-series variations simultaneously for better time series forecasting, which is implemented by channel-wise self-attention and autoregressive self-attention. To evaluate the effectiveness of our method, we conducted experiments across eight open-source datasets and compared it with the state-of-the-art methods. Through the comparison results, our Leddam (LEarnable Decomposition and Dual Attention Module) not only demonstrates significant advancements in predictive performance but also the proposed decomposition strategy can be plugged into other methods with a large performance-boosting, from 11.87% to 48.56% MSE error degradation.
@@ 1 Introduction
The rising demands in diverse real-world domains have generated an urgent requirement for precise multivariate time series forecasting methodologies, as demonstrated in fields like energy management (Dong et al., 2023; Liu et al., 2023b; Yi et al., 2023a), weather forecasting (Anonymous, 2024b, c, a), disease control (Yi et al., 2023b; Liu et al., 2023a; Zhou et al., 2022b; Ni et al., 2023), and traffic planning (Rangapuram et al., 2018; Zhao et al., 2017; Shao et al., 2022). The foundation of precise forecasting models lies in effectively identifying and modelling intricate patterns embedded in multivariate time series. Two primary patterns (Figure 1(a)) emerge as inter-series dependencies and intra-series variations (Zhang & Yan, 2023). The former delineates the intricate interplay and correlations among distinct variables, while the latter encapsulates both enduring and ephemeral fluctuations within each specific time series.
@@
Figure 1: (a) Demonstration of inter-series dependencies and intra-series variations. (b) Visualization of different decomposition schemes in Electricity data. RAW means the raw time series. MOV means Moving average kernel, LD means our learnable decomposition module.
@@
However, the time series of each constituent variable in multivariate time series data often displays distinctive variations in its trends. Such discrepancies inherent in raw time series may complicate the modelling of inter-series dependencies (Liu et al., 2024). Furthermore, time series in the real world are continually susceptible to distributional shifts induced by the evolution of their trends—a distinctive attribute that adds complexity to modelling the dynamic intra-series variation patterns within the sequences (Taylor & Letham, 2018; Liu et al., 2022b). Therefore, a robust forecasting method should be capable of addressing the following two challenges: (1) How to precisely unravel patterns within raw time series under the interference of trend components. (2) How to efficiently model inter-series dependencies and intra-series variations.
@@
Certain studies (Wu et al., 2021; Zhou et al., 2022c; Wang et al., 2023; Zeng et al., 2023) aim to address the first challenge by conducting trend-seasonal decomposition of the original time series using moving average kernels (MOV). However, such an untrainable procedure along with moving average kernels leads to a lack of robustness. Moreover, the uniform assignment of weights to each data point within the sliding window may impede their ability to discern specific patterns. Such limitation becomes apparent when dealing with intricate time series data (RAW data), particularly those with non-linear structures or significant noise levels, as illustrated in Figure 1(b) (RAW VS. MOV). Therefore, it is necessary to develop a learnable decomposition method to revitalize multivariate time series forecasting tasks.
@@
For the second challenge of modelling inter-series dependencies and intra-series variations in time series forecasting, recent efforts have turned to Transformer architectures renowned for their robust pairwise dependency delineation and multi-level representation extraction within sequences (Vaswani et al., 2017). iTransformer (Liu et al., 2024) effectively handles inter-series dependencies through ‘Channel-wise self-attention’, embedding entire time series into a token but lacks explicit learning of intra-series variations. Other works directly employ pair-wise attention mechanisms for intra-series variations (Liu et al., 2022a; Zhou et al., 2022a; Li et al., 2019; Liu et al., 2022b). However, they improperly use permutation-invariant attention mechanisms on the temporal dimension (Zeng et al., 2023). Alternatively, some approaches shift to partitioning time series into patches and applying self-attention modelling on these patches (Nie et al., 2023; Zhang & Yan, 2023). Yet, such methods inherently lead to information loss as patches encapsulate only a portion of the original sequence. Moreover, the optimal patch length is also hard to determine. Therefore, we aim to deal with it by generating features suitable for modelling intra-series variations while maximizing information preservation to avoid the above disadvantages.
@@ 2 Related work
Time Series Data Decomposition. Due to the capacity of the moving average kernel to extract global information from sequence data, Autoformer (Wu et al., 2021) initially proposed employing it for extracting the trend part of time series. Later works, including MICN (Wang et al., 2023), FEDformer (Zhou et al., 2022c), DLinear (Zeng et al., 2023), etc., have predominantly adhered to their methodology. However, a rudimentary averaging kernel may inadequately capture precise trends in time series characterized by more intricate patterns than simple linear relationships.
@@
Intra-series variations modeling. To a precise time series forecasting model, comprehending the temporal variations inherent within the time series is of paramount importance. Due to the intrinsic limitations of point-to-point attention mechanisms, such as those employed in models like Informer (Zhou et al., 2022a), Reformer (Kitaev et al., 2020), and Pyraformer (Liu et al., 2022a), the resultant attention maps are prone to suboptimality. This arises from the fact that individual points in a time series, in contrast to words (Vaswani et al., 2017) or image patches (Dosovitskiy et al., 2021), lack explicit semantic information. Subsequent endeavors involve partitioning the primary time series into a series of patches (Zhang & Yan, 2023; Nie et al., 2023), followed by the application of self-attention mechanisms across these patches to model temporal variations. However, segmenting time series into patches inevitably introduces information loss.
@@
Inter-series dependencies modeling. Inter-series dependencies constitute a pivotal attribute that distinguishes multivariate time series from their univariate counterparts. It has come to our attention that the majority of transformer-based methodologies opt to treat values from different variables at the same time step or from distinct channels as tokens (Zhou et al., 2022a; Kitaev et al., 2020; Liu et al., 2022a, b) to model inter-series dependencies. Such strategies may result in attention maps that lack meaningful information, consequently impeding the effective and accurate modelling of the information we seek (Liu et al., 2024). Certain endeavours have sought to adopt channel-independent designs, aiming to mitigate the reduction in predictive accuracy induced by this operation, as exemplified by PatchTST (Nie et al., 2023) and DLinear (Zeng et al., 2023). Channel Independence (CI) regarding variates of time series independently and adopting the shared backbone, have gained increasing popularity in forecasting with performance promotions as an architecture-free method. Recent works (Han et al., 2023; Li et al., 2023) found that while Channel Dependence (CD) benefits from a higher capacity ideally, CI can greatly boost the performance because of sample scarcity, since most of the current forecasting benchmarks are not large enough. However, neglecting the interdependencies among variables may lead to suboptimal outcomes (Liu et al., 2024). Crossformer (Zhang & Yan, 2023) strives to partition time series into patches and subsequently engage in the learning of inter-series dependencies on these patches, a process which, as argued previously, may engender information loss. iTranformer (Liu et al., 2024) achieves more precise modelling of relationships among variables of multivariate time series by embedding the entire time series of a variate into a token, thereby avoiding information loss.
@@ 3.1 Problem Definition
3.2 Learnable Convolutional Decomposition
We employ a superior learnable 1D convolutional decomposition kernel instead of moving average kernel to comprehensively encapsulate the nuanced temporal variations in the time series.
Figure 3: General process of ‘Dual Attention Module’ to deal with Inter-series dependencies and Intra-series variations respectively. ‘Channel-wise self-attention’ embeds the whole series of a channel to generate ‘Whole Series Embedding’, and transformer encoders are employed to model Inter-series dependencies. ‘Auto-regressive self-attention’ generates ‘Auto-regressive Embedding’ and still utilizes transformer encoders to model Intra-series variations
Projection and position embedding. Following iTransformer (Liu et al., 2024), we first map time series data $X\in\mathbb{R}^{N\times T}$ from the original space to a new space, subsequently incorporating positional encoding $Pos\in\mathbb{R}^{N\times D}$ as $X_{embed}\in\mathbb{R}^{N\times D}$ following $X_{embed}=(XW+b)+Pos,$
with weights $W\in\mathbb{R}^{T\times D},b\in\mathbb{R}^{1\times D}$, where $D$ is the dimension of the layer.
@@ 3.2 Learnable Convolutional Decomposition
Trend part. Given the smoother and more predicatblenature of the trend part, we employ a simple MLP for projection to derive the trend part’s output akin to (Zeng et al., 2023; Wang et al., 2023), which reads:
@@
Seasonal part. Considering the suitability of the seasonal component for modelling inter-series dependencies and intra-series variations, we transform $X_{Seasonal}\in\mathbb{R}^{N\times D}$ into two distinct embeddings: ‘Whole Series Embedding’ and ‘Auto-regressive Embedding’. This facilitates the modelling and learning processes for the two patterns.
@@ 3.3 Dual Attention Module
We propose a ‘Dual Attention Module’ to model the inter-series dependencies and intra-series variations simultaneously. Concretely, we devise ‘Channel-wise self-attention’ for modelling the former and ‘Auto-regressive self-attention’ for modelling the latter.
Inter-series dependencies modelling. To model inter-series dependencies, follow iTranformer (Liu et al., 2024), we consider $X_{Seasonal}[i,:]\in\mathbb{R}^{1\times D},\quad i=1,2,\ldots,N$ as a token as shown in Figure 3. Subsequently, all tokens are sent into a vanilla transformer encoder for learning purposes to get $X_{Inter}\in\mathbb{R}^{N\times D}$:
@@
We denote this process as ‘Channel-wise self-attention’ and the generated embeddings as ‘Whole Series Embedding’, which, while, maintaining most information of the sequence in contrast to patches or segments, is better suitable for modelling inter-series dependencies, as all semantic information of the variates are saved.
Intra-series variations modelling. To capture intra-series variations, we propose an advanced methodology. Herein, the initial sequence undergoes auto-regressive processing, generating tokens that, while meticulously retaining the entirety of the original information, partially emulate the dynamic variations present in raw time series.
@@ 4.1 Experimental Settings
In this section, we first introduce the whole experiment settings under a fair comparison. Secondly, we illustrate the experiment results by comparing Leddam with the 8 current state-of-the-art (SOTA) methods. Further, we conducted an ablation study to comprehensively investigate the effectiveness of the learnable convolutional decomposition module and the effectiveness of the introduced ‘Dual attention module’.
@@
ETTh1 0.428 0.428 0.454 0.447 0.458 0.450 0.561 0.535 0.456 0.452 0.449 0.442 0.624 0.575 0.434 0.437 0.486 0.467
ETTh2 0.361 0.391 0.383 0.407 0.414 0.427 0.587 0.525 0.559 0.515 0.387 0.407 0.942 0.684 0.611 0.550 0.954 0.723
ETTm1 0.385 0.397 0.407 0.410 0.400 0.406 0.392 0.414 0.403 0.407 0.387 0.400 0.513 0.509 0.403 0.427 0.411 0.418
ETTm2 0.280 0.325 0.288 0.332 0.291 0.333 0.328 0.382 0.350 0.401 0.283 0.327 1.219 0.827 0.293 0.336 0.310 0.347
Electricity 0.165 0.260 0.178 0.270 0.192 0.295 0.187 0.295 0.212 0.300 0.216 0.304 0.244 0.334 0.251 0.344 0.268 0.365
Solar-Energy 0.228 0.262 0.233 0.262 0.301 0.319 0.296 0.371 0.330 0.401 0.270 0.307 0.641 0.639 0.347 0.417 0.282 0.375
Traffic 0.456 0.286 0.428 0.282 0.620 0.336 0.542 0.315 0.625 0.383 0.555 0.362 0.550 0.304 0.760 0.473 0.804 0.509
Weather 0.241 0.271 0.258 0.279 0.259 0.287 0.243 0.299 0.265 0.317 0.259 0.281 0.259 0.315 0.271 0.320 0.292 0.363
$1^{st}$ Count 7 7 1 1 0 0 0 0 0 0 0 0 0 0 0 0 0 0
@@ 4.2 Experiments Results
Quantitative comparison. Comprehensive forecasting results are listed in Table 2 with the best bold in red and the second underlined in blue. We leave full forecasting results in APPENDIX E to save place. The lower MSE/MAE indicates the better prediction result. It is unequivocally evident that Leddam has demonstrated superior predictive performance across all datasets except Traffic, in which iTransformer gets the best forecasting performance. Similar to PatchTST and iTransformer, Leddam employs a vanilla transformer encoder as its backbone, devoid of any structural modifications. It is noteworthy, however, that these three models consistently exhibit superior performance across all datasets. This at least partially indicates that, compared to intricately designed model architectures, superior representation of raw time series features, such as ‘Whole Series Embedding’ used in iTransformer and Leddam along with PatchTST’s patches, may indeed constitute the pivotal factors for achieving more efficient time-series predictions. It is noteworthy that among these three models, PacthTST employs a channel-independent design, exclusively addressing intra-series variations without giving consideration to inter-series dependencies. iTransformer utilizes channel-wise self-attention to model inter-series dependencies but falls short of adequately capturing intra-series variations. In comparison to both, the proposed Leddam incorporates ‘Whole Series Embedding’ and ‘Auto-regressive Embedding’ to respectively model inter-series dependencies and intra-series variations. Consequently, Leddam demonstrates superior performance across various datasets. However, these three models still maintain a leading position over other models across most datasets. This aligns with our hypothesis that appropriately modelling inter-series dependencies and intra-series variations in multivariate time series is key for achieving more precise forecasting.

### v5

@@ Abstract
Predicting multivariate time series is crucial, demanding precise modeling of intricate patterns, including inter-series dependencies and intra-series variations. Distinctive trend characteristics in each time series pose challenges, and existing methods, relying on basic moving average kernels, may struggle with the non-linear structure and complex trends in real-world data. Given that, we introduce a learnable decomposition strategy to capture dynamic trend information more reasonably. Additionally, we propose a dual attention module tailored to capture inter-series dependencies and intra-series variations simultaneously for better time series forecasting, which is implemented by channel-wise self-attention and autoregressive self-attention. To evaluate the effectiveness of our method, we conducted experiments across eight open-source datasets and compared it with the state-of-the-art methods. Through the comparison results, our Leddam (LEarnable Decomposition and Dual Attention Module) not only demonstrates significant advancements in predictive performance but also the proposed decomposition strategy can be plugged into other methods with a large performance-boosting, from 11.87% to 48.56% MSE error degradation. Code is available at this link: https://github.com/Levi-Ackman/Leddam.
@@ 1 Introduction
The rising demands in diverse real-world domains have generated an urgent requirement for precise multivariate time series forecasting methodologies, as demonstrated in fields like energy management (Dong et al., 2023; Wu et al., 2024; Yi et al., 2023a), weather forecasting (Anonymous, 2024b, c, a), disease control (Yi et al., 2023b; Liu et al., 2023; Zhou et al., 2022b; Ni et al., 2023), and traffic planning (Rangapuram et al., 2018; Zhao et al., 2017; Shao et al., 2022). The foundation of precise forecasting models lies in effectively identifying and modeling intricate patterns embedded in multivariate time series. Two primary patterns (Figure 1(a)) emerge as inter-series dependencies and intra-series variations (Zhang & Yan, 2023). The former delineates the intricate interplay and correlations among distinct variables, while the latter encapsulates both enduring and ephemeral fluctuations within each specific time series.
@@
Figure 1: (a) Demonstration of inter-series dependencies and intra-series variations. (b) Visualization of different decomposition schemes in Electricity data. RAW means the raw time series. MOV means moving average kernel, and LD means our learnable decomposition module.
@@
However, the time series of each constituent variable in multivariate time series data often displays distinctive variations in its trends. Such discrepancies inherent in raw time series may complicate the modeling of inter-series dependencies (Liu et al., 2024). Furthermore, time series in the real world are continually susceptible to distributional shifts induced by the evolution of their trends—a distinctive attribute that adds complexity to modeling the dynamic intra-series variation patterns within the sequences (Taylor & Letham, 2018; Liu et al., 2022b). Therefore, a robust forecasting method should be capable of addressing the following two challenges: (1) How to precisely unravel patterns within raw time series under the interference of trend components. (2) How to efficiently model inter-series dependencies and intra-series variations.
@@
Certain studies (Wu et al., 2021; Zhou et al., 2022c; Wang et al., 2023; Zeng et al., 2023) aim to address the first challenge by conducting trend-seasonal decomposition of the original time series using Moving Average kernel (MOV). However, such an untrainable procedure along with moving average kernels leads to a lack of robustness. Moreover, the uniform assignment of weights to each data point within the sliding window may impede their ability to discern specific patterns. Such limitation becomes apparent when dealing with intricate time series data (RAW data), particularly those with non-linear structures or significant noise levels, as illustrated in Figure 1(b) (RAW VS. MOV). Therefore, it is necessary to develop a learnable decomposition method to revitalize multivariate time series forecasting tasks.
@@
For the second challenge of modeling inter-series dependencies and intra-series variations in time series forecasting, recent efforts have turned to Transformer architectures renowned for their robust pairwise dependency delineation and multi-level representation extraction within sequences (Vaswani et al., 2017). iTransformer (Liu et al., 2024) effectively handles inter-series dependencies through ‘Channel-wise self-attention’, embedding entire time series into a token but lacks explicit learning of intra-series variations. Other works directly employ pair-wise attention mechanisms for intra-series variations (Liu et al., 2022a; Zhou et al., 2022a; Li et al., 2019; Liu et al., 2022b). However, they improperly use permutation-invariant attention mechanisms on the temporal dimension (Zeng et al., 2023). Alternatively, some approaches shift to partitioning time series into patches and applying self-attention modeling on these patches (Nie et al., 2023; Zhang & Yan, 2023). Yet, such methods inherently lead to information loss as patches encapsulate only a portion of the original sequence. Moreover, the optimal patch length is also hard to determine. Therefore, we aim to deal with it by generating features suitable for modeling intra-series variations while maximizing information preservation to avoid the above disadvantages.
@@ 2 Related work
Time Series Data Decomposition. Due to the capacity of the moving average kernel to smooth out short-term fluctuations or noise in the time series, Autoformer (Wu et al., 2021) initially proposed employing the moving average kernel for extracting the trend part of the time series. Later works, including MICN (Wang et al., 2023), FEDformer (Zhou et al., 2022c), DLinear (Zeng et al., 2023), etc., have predominantly adhered to their methodology. However, a rudimentary averaging kernel may inadequately capture precise trends in time series characterized by more intricate patterns than simple linear relationships. It applies uniform weighting to all data points within the window size, which may not be suitable for capturing non-linear or non-stationary trends present in the data.
@@
Intra-series Variations Modeling. It is important to comprehend the temporal variations inherent within the time series for a precise time series forecasting model. Due to the intrinsic limitations of point-to-point attention mechanisms, such as those employed in models like Informer (Zhou et al., 2022a), Reformer (Kitaev et al., 2020), and Pyraformer (Liu et al., 2022a), the resultant attention maps are prone to suboptimality. This arises from the fact that individual points in a time series, in contrast to words (Vaswani et al., 2017) or image patches (Dosovitskiy et al., 2021), lack explicit semantic information. Subsequent endeavours involve partitioning the primary time series into a series of patches (Zhang & Yan, 2023; Nie et al., 2023), followed by applying self-attention mechanisms across these patches to model temporal variations. However, segmenting time series into patches inevitably introduces information loss.
@@
Inter-series Dependencies Modeling. Inter-series dependencies constitute a pivotal attribute that distinguishes multivariate time series from their univariate counterparts. It has come to our attention that the majority of transformer-based methodologies opt to treat values from different variables at the same time step or from distinct channels as tokens (Zhou et al., 2022a; Kitaev et al., 2020; Liu et al., 2022a, b) to model inter-series dependencies. Such strategies may result in attention maps that lack meaningful information, consequently impeding the effective and accurate modeling of the information we seek (Liu et al., 2024). Certain endeavours have sought to adopt channel-independent designs, aiming to mitigate the reduction in predictive accuracy induced by this operation, as exemplified by PatchTST (Nie et al., 2023) and DLinear (Zeng et al., 2023). Channel Independence (CI) regarding variates of time series independently and adopting the shared backbone, have gained consistently increasing popularity in forecasting with performance promotions as an architecture-free method. Recent works (Han et al., 2023; Li et al., 2023) found that while Channel Dependence (CD) benefits from a higher capacity ideally, CI can greatly boost the performance because of sample scarcity, since most of the current forecasting benchmarks are not large enough. However, neglecting the interdependencies among variables may lead to suboptimal outcomes (Liu et al., 2024). Crossformer (Zhang & Yan, 2023) strives to partition time series into patches and subsequently engage in the learning of inter-series dependencies on these patches, a process which, as argued previously, may engender information loss. iTranformer (Liu et al., 2024) achieves more precise modeling of relationships among variables of multivariate time series by embedding the entire time series of a variate into a token, thereby avoiding information loss.
@@ 3.1 Problem Definition
3.2 Learnable Decomposition Module
We employ a superior learnable 1D convolutional decomposition kernel instead of a moving average kernel to comprehensively encapsulate the nuanced temporal variations in the time series.
Figure 3: General process of ‘Dual Attention Module’ to deal with Inter-series dependencies and Intra-series variations, respectively. ‘Channel-wise self-attention’ embeds the whole series of a channel to generate ‘Whole Series Embedding’, and transformer encoders are employed to model Inter-series dependencies. ‘Auto-regressive self-attention’ generates ‘Auto-regressive Embedding’ and still utilizes transformer encoders to model Intra-series variations
Projection and position embedding. Following iTransformer (Liu et al., 2024), we first map time series data $X\in\mathbb{R}^{N\times T}$ from the original space to a new space, subsequently incorporating positional encoding $Pos\in\mathbb{R}^{N\times D}$ as $X_{embed}\in\mathbb{R}^{N\times D}$ following $X_{embed}=(XW+b)+Pos,$ with weights $W\in\mathbb{R}^{T\times D},b\in\mathbb{R}^{1\times D}$, where $D$ is the dimension of the layer.
@@ 3.2 Learnable Decomposition Module
Trend part. Given the smoother and more predictable nature of the trend part, we employ a simple MLP for projection to derive the trend part’s output akin to (Zeng et al., 2023; Wang et al., 2023), which reads:
@@
Seasonal part. Considering the suitability of the seasonal component for modeling inter-series dependencies and intra-series variations, we transform $X_{Seasonal}\in\mathbb{R}^{N\times D}$ into two distinct embeddings: ‘Whole Series Embedding’ and ‘Auto-regressive Embedding’. This facilitates the modeling and learning processes for the two patterns.
@@ 3.3 Dual Attention Module
We propose a ‘Dual Attention Module’ to model the inter-series dependencies and intra-series variations simultaneously. Concretely, we devise ‘Channel-wise self-attention’ for modeling the former and ‘Auto-regressive self-attention’ for modeling the latter.
Inter-series dependencies modeling. To model inter-series dependencies, follow iTranformer (Liu et al., 2024), we consider $X_{Seasonal}[i,:]\in\mathbb{R}^{1\times D},\quad i=1,2,\ldots,N$ as a token as shown in Figure 3. Subsequently, all tokens are sent into a vanilla transformer encoder for learning purposes to get $X_{Inter}\in\mathbb{R}^{N\times D}$:
@@
We denote this process as ‘Channel-wise self-attention’ and the generated embeddings as ‘Whole Series Embedding’, which, while, maintaining most information of the sequence in contrast to patches or segments, is better suitable for modeling inter-series dependencies, as all semantic information of the variates are saved.
Intra-series variations modeling. To capture intra-series variations, we propose an advanced methodology. Herein, the initial sequence undergoes auto-regressive processing, generating tokens that, while meticulously retaining the entirety of the original information, partially emulate the dynamic variations present in raw time series.
@@ 4.1 Experimental Settings
In this section, we first introduce the whole experiment settings under a fair comparison. Secondly, we illustrate the experiment results by comparing Leddam with the eight current state-of-the-art (SOTA) methods. Further, we conducted an ablation study to comprehensively investigate the effectiveness of the learnable convolutional decomposition module and the effectiveness of the Dual attention module.
@@
ETTh1 0.431 0.429 0.454 0.447 0.458 0.450 0.561 0.535 0.456 0.452 0.449 0.442 0.624 0.575 0.434 0.437 0.486 0.467
ETTh2 0.373 0.399 0.383 0.407 0.414 0.427 0.587 0.525 0.559 0.515 0.387 0.407 0.942 0.684 0.611 0.550 0.954 0.723
ETTm1 0.386 0.397 0.407 0.410 0.400 0.406 0.392 0.414 0.403 0.407 0.387 0.400 0.513 0.509 0.403 0.427 0.411 0.418
ETTm2 0.281 0.325 0.288 0.332 0.291 0.333 0.328 0.382 0.350 0.401 0.283 0.327 1.219 0.827 0.293 0.336 0.310 0.347
Electricity 0.169 0.263 0.178 0.270 0.192 0.295 0.187 0.295 0.212 0.300 0.216 0.304 0.244 0.334 0.251 0.344 0.268 0.365
Solar-Energy 0.230 0.264 0.233 0.262 0.301 0.319 0.296 0.371 0.330 0.401 0.270 0.307 0.641 0.639 0.347 0.417 0.282 0.375
Traffic 0.467 0.294 0.428 0.282 0.620 0.336 0.542 0.315 0.625 0.383 0.555 0.362 0.550 0.304 0.760 0.473 0.804 0.509
Weather 0.242 0.272 0.258 0.279 0.259 0.287 0.243 0.299 0.265 0.317 0.259 0.281 0.259 0.315 0.271 0.320 0.292 0.363
$1^{st}$ Count 7 6 1 2 0 0 0 0 0 0 0 0 0 0 0 0 0 0
@@ 4.2 Experiments Results
Quantitative comparison. Comprehensive forecasting results are listed in Table 2 with the best bold in red and the second underlined in blue. We leave full forecasting results in APPENDIX G to save place. The lower MSE/MAE indicates the better prediction result. It is unequivocally evident that Leddam has demonstrated superior predictive performance across all datasets except Traffic, in which iTransformer gets the best forecasting performance. Like PatchTST and iTransformer, Leddam employs a vanilla transformer encoder as its backbone, devoid of any structural modifications. It is noteworthy, however, that these three models consistently exhibit superior performance across all datasets. This at least partially indicates that, compared to intricately designed model architectures, superior representation of raw time series features, such as ‘Whole Series Embedding’ used in iTransformer and Leddam along with PatchTST’s patches, may indeed constitute the pivotal factors for achieving more efficient time-series predictions. It is noteworthy that among these three models, PacthTST employs a channel-independent design, exclusively addressing intra-series variations without considering inter-series dependencies. iTransformer utilizes channel-wise self-attention to model inter-series dependencies but falls short of adequately capturing intra-series variations. Compared to both, the proposed Leddam incorporates ‘Whole Series Embedding’ and ‘Auto-regressive Embedding’ to model inter-series dependencies and intra-series variations. Consequently, Leddam demonstrates superior performance across various datasets. However, these three models still maintain a leading position over other models across most datasets. This aligns with our hypothesis that appropriately modeling inter-series dependencies and intra-series variations in multivariate time series is key for achieving more precise forecasting.